In a Poisson distribution, μ=.30μ=.30. (Round your answers to 4 decimal places.)
  
(a) What is the probability that x=0x=0?
  
  Probability   
(b) What is the probability that x>1x>1?
  
  Probability   
Answer #1

X ~ poi ()

Where = 0.30

Poisson probability distribution is

P(X) = e- * x / X!

a)

P(X = 0) = e-0.30

= 0.7408

b)

P(X > 1) = 1 - P( X <= 1)

= 1 - [ P(X = 0) + P(X = 1) ]

= 1 - [ e-0.30 + e-0.30 * 0.30 ]

= 0.0369